# FaCT: Faithful Concept Traces for Explaining Neural Network Decisions

### v1

@@ Abstract
[…]
@@ 1 Introduction
To address this, concept-based explanation methods have become a popular tool to decompose model decisions [9, 34] and arbitrary activations [17, 21] into high-level human-interpretable concepts. These include part-prototype networks [9, 55] and concept bottleneck models [34, 40, 47] that aim to create inherently interpretable models, yet often provide unfaithful explanations [26, 37, 57]. Other approaches like CRAFT [17] decompose model activations post hoc, but the extracted concepts are not directly used for prediction, causing reliance on approximate methods to estimate the importance of concepts to the output [16] and to visualize which region of the input the concept is activated for [50, 57, 59], which may also not be faithful to the original decision-making [2]. Furthermore, the concepts are often subject to restrictive assumptions, e.g. being class-specific [17, 21, 35], being limited to small patches or object parts [17, 21, 55], or belonging to a pre-defined set [34, 40, 47], hindering the methods from faithfully explaining the concepts used by the model.
@@
Since the concepts used by the model may not align with any predefined human-annotated object parts (see Fig.˜3, where annotations fail to capture our concepts), we also find that existing metrics for evaluating concept consistency [30] that make such assumptions are suboptimal. To address this, inspired by its recent success for semantic correspondence [3, 62, 63], we utilize DINOv2 [42] features to evaluate concept consistency without human annotations. Our proposed C2-score takes into account the input features that activate the concept and evaluates consistency independent of pre-defined annotations, while correlating well with human notion of consistency.
@@
Our models remain competitive on ImageNet, while providing faithful concept-based explanation with a diverse (shared) concept basis. We quantitatively demonstrate how our concepts are more consistent, and through a user-study, how they are more interpretable than baselines.
@@ 2 Related Work
Evaluating Explanation Methods is essential to ensure that explanations can be trusted and are interpretable. For attribution methods, many works proposed sanity checks for measuring their faithfulness [2, 33], metrics for evaluating their localization ability [51, 56], as well as synthetic datasets [25, 48] for a controlled evaluation. For concept-based explanations, prior work [30] proposes evaluating concepts in terms of binarized attributions having high IoU with part annotations. This however is limited to class-specific concepts, requires annotations for every part, and assumes every concept should correspond to an annotated part. In this work, we propose a novel consistency metric C2-score, which leverages foundation models to evaluate consistency of concepts in a class-agnostic manner. Closer to our metric are works that measure ’monosemanticity’ of individual neurons [15] as the similarity of set of images or crops. In contrast, our work considers full-image attribution and, through appropriate baselines, can evaluate different concept extraction methods.
@@ 3 Faithfully Explaining with Concepts
Sparsity can be induced by e.g. $\ell_{1}$ regularization on the sparse codes, or by explicitly enforcing k-sparse codes, i.e. TopK-SAE [20]. In our work, we use TopK-SAEs, with the encoder having no biases, so the sparse codes can be faithfully represented as a strictly linear combination of features.
@@ 4 Concept Consistency Evaluation
Therefore, C2-score lends itself as a simple yet robust evaluation framework that takes the concept attribution into account and can evaluate both class-specific and shared concept sets. Having such an evaluation framework, in Section˜5.2 we compare our concepts with prior work. In Fig.˜4-right, we further validate the proposed C2-score and how it is close to human definition of consistency, by showing randomly sampled concepts with high and low C2-score. In Sec.˜C we further elaborate on C2-score, showing how it outperforms annotations and how it correlates with user interpretability.
@@ 5 Inspecting the Concepts
Before we begin, let us briefly discuss our implementation details. We used ImageNet-pretrained B-cos [7] networks with fixed parameters and used bias-free TopK-SAE [20] sweeping $TopK\in(8,16,32)$, with $K\in(8192,16384)$ total concepts. We constructed the training dataset by sampling features patches from ImageNet’s training set. Throughout the paper, each concept is visualized by top-activating test images, with attribution (Eq.˜12) and the image category. See also Sec.˜A.
@@ Acknowledgments
We would like to thank our colleagues Thomas Wimmer, Olaf Dünkel, Moritz Böhle, Nhi Pham, and Siddhartha Gairola for the helpful discussions. We are especially thankful to Thomas for helping with Fig.˜1.
@@ Appendix A Training Details
For the SAEs, we base our implementation on the Dictionary-Learning codebase [26]. We use the TopK-SAE [14], and modify it to be bias-free (i.e. no data whitening and no biases for encoder or decoder). This is needed so that faithful output contribution in Eq.˜9 and faithful input attribution in Eq.˜12 can be obtained.

### v2

@@ Abstract
Deep networks have shown remarkable performance across a wide range of tasks, yet getting a global concept-level understanding of how they function remains a key challenge. Many post-hoc concept-based approaches have been introduced to understand their workings, yet they are not always faithful to the model. Further, they make restrictive assumptions on the concepts a model learns, such as class-specificity, small spatial extent, or alignment to human expectations. In this work, we put emphasis on the faithfulness of such concept-based explanations and propose a new model with model-inherent mechanistic concept-explanations. Our concepts are shared across classes and, from any layer, their contribution to the logit and their input-visualization can be faithfully traced. We also leverage foundation models to propose a new concept-consistency metric, C2-score, that can be used to evaluate concept-based methods. Compared to prior work, we show that our concepts are quantitatively more consistent and that users find them to be more interpretable, while retaining competitive ImageNet performance.  [^1] Code available at github.com/m-parchami/FaCT. Denotes equal contribution as advisors.
@@ 1 Introduction
To address this, concept-based explanation methods have become a popular tool to decompose model decisions [10, 41] and arbitrary activations [22, 26] into high-level human-interpretable concepts. These include part-prototype networks [10, 66] and concept bottleneck models [41, 49, 56] that aim to create inherently interpretable models, yet often provide unfaithful explanations [31, 45, 72]. Other approaches like CRAFT [22] decompose model activations post hoc, but the extracted concepts are not directly used for prediction, leading to reliance on approximate methods to estimate the importance of concepts to the output [19] and to visualize which region of the input the concept is activated for [59, 72, 74], which may also not be faithful to the original decision-making [2]. Furthermore, concepts are often subject to restrictive assumptions, e.g. being class-specific [22, 26, 42], being limited to small patches or object parts [22, 26, 66], or belonging to a pre-defined set [41, 49], hindering such methods from faithfully explaining the true concepts used by the model.
@@
Since the concepts used by the model may not align with any predefined human-annotated object parts (see Fig.˜3, where annotations fail to capture our concepts), we also find that existing metrics for evaluating concept consistency [35] that make such assumptions are suboptimal. To address this, inspired by its recent success for semantic correspondence [3, 77, 78], we utilize DINOv2 [51] features to evaluate concept consistency without human annotations. Our proposed C2-score takes into account the input features that activate the concept and evaluates consistency independent of pre-defined annotations, while correlating well with human notions of consistency.
@@
Our models remain competitive on ImageNet [14], while providing faithful concept-based explanations with a diverse (shared) concept basis. We quantitatively demonstrate that our concepts are more consistent, and through a user-study, that they are more interpretable than baselines.
@@ 2 Related Work
Evaluating Explanation Methods is essential to ensure that explanations can be trusted and are interpretable. For attribution methods, many works proposed sanity checks for measuring their faithfulness [2, 39], metrics for evaluating their localization ability [60, 68], as well as synthetic datasets [30, 57] for a controlled evaluation. For concept-based explanations, prior work [35] proposes evaluating concepts in terms of binarized attributions having high IoU with part annotations. This however is limited to class-specific concepts, requires annotations for every part, and assumes every concept should correspond to an annotated part. In this work, we propose a novel consistency metric C2-score, which leverages foundation models to evaluate consistency of concepts in a class-agnostic manner. Close to our metric are works that measure ‘monosemanticity’ of individual neurons [18] as the similarity of a set of images or crops. In contrast, our work considers full-image attributions and, through appropriate baselines, can evaluate various concept extraction methods.
@@ 3 Faithfully Explaining with Concepts
Sparsity can be induced by e.g. $\ell_{1}$ regularization on the sparse codes, or by explicitly enforcing k-sparse codes, i.e. TopK-SAE [25, 44]. In our work, we use TopK-SAEs, with the encoder having no biases, so the sparse codes can be faithfully represented as a strictly linear combination of features.
@@ 4 Concept Consistency Evaluation
Therefore, C2-score lends itself as a simple yet robust evaluation framework that takes the concept attribution into account and can evaluate both class-specific and shared concept sets. Having such an evaluation framework, in Section˜5.2 we compare our concepts with prior work. In Fig.˜4-right, we further validate the proposed C2-score and how it is close to human definition of consistency, by showing randomly sampled concepts with high and low C2-score. In Sec.˜C we elaborate on C2-score, showing how it outperforms annotations and how it correlates with user interpretability.
@@ 5 Inspecting the Concepts
Before we begin, let us briefly discuss our implementation details. We used ImageNet-pretrained B-cos [8] networks with fixed parameters and used bias-free TopK-SAE [25, 44] sweeping $TopK\in(8,16,32)$, with $K\in(8192,16384)$ total concepts. We constructed the training dataset by sampling feature patches from ImageNet’s training set. Throughout the paper, each concept is visualized by top-activating test images, with attribution (Eq.˜12) and the image category. See also Sec.˜A.
@@ Acknowledgments
Sukrut Rao and Bernt Schiele were funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – GRK 2853/1 “Neuroexplicit Models of Language, Vision, and Action” - project number 471607914. We would like to thank our colleagues Thomas Wimmer, Olaf Dünkel, Moritz Böhle, Nhi Pham, and Siddhartha Gairola for helpful discussions. We are especially thankful to Thomas for helping with Fig.˜1.
@@ Appendix A Training Details
For the SAEs, we base our implementation on the Dictionary-Learning codebase [46]. We use the TopK-SAE [25, 44], and modify it to be bias-free (i.e. no data whitening and no biases for encoder or decoder). This is needed so that faithful output contribution in Eq.˜9 and faithful input attribution in Eq.˜12 can be obtained.